# The (110) surface is the key to the high ice nucleation efficiency of potassium feldspar

### v1

@@ Abstract
Mineral dust aerosols strongly influence Earth’s climate by acting as ice-nucleating particles (INPs). Feldspar minerals, particularly K-feldspar, are recognized as dominant INPs, and a previous study attributed this behavior to (100) surfaces exposed at defects. Using machine-learning molecular dynamics simulations, we systematically investigate ice nucleation on all K-feldspar surfaces. We identify the (110) surface, exposed at defects such as steps, as the most active plane for ice formation. This surface uniquely structures interfacial water into an arrangement resembling that on the (110) surface of cubic ice, providing an optimal template for nucleation. Using advanced sampling, we directly observe the formation of clusters with cubic-ice structure and their orientation agrees with experiment. These results provide a microscopic explanation of how ice forms in our planet’s atmosphere.
@@ Introduction
Figure 1: K-feldspar serves as an effective ice-nucleating particle in the atmosphere owing to structural defects that promote ice nucleation. (A) Schematic illustration of the heterogeneous ice nucleation at K-feldspar defects with exposed non-perfect cleavage planes. The dotted and dashed rectangles represent the first and second prismatic surfaces of ice, which are approximately parallel to the (100) and (110) planes of feldspar, respectively. (B) The triclinic crystal structure of microcline K-feldspar. (C) Crystallographic planes of K-feldspar: (001), (010), (100), (110), ($\bar{1}$10), and ($\bar{2}$01). (D) Configuration for molecular dynamics simulations with one termination of the K-feldspar (110) surface in contact with water above.
In this work, we use MD simulations to study a broad range of K-feldspar surfaces exposed to water, including the (001), (010), (100), (110), ($\bar{1}$10), and ($\bar{2}$01) surfaces. The interatomic forces driving these simulations are derived from a machine-learning interatomic potential (MLP) trained to achieve a very high accuracy using quantum-mechanical electronic-structure theory calculations. Our simulations reveal that the K-feldspar (110) surface structures interfacial liquid water into an arrangement closely resembling that on the (110) surface of cubic ice (equivalent to the secondary prismatic surface of hexagonal ice). In contrast, water at all other feldspar surfaces does not show signatures of ice-like structure. We also carried out enhanced sampling molecular simulations in order to bring ice nucleation, which is a rare event in the timescale of standard MD simulations, within reach. The ice clusters that nucleate at the (110) surface during these simulations exhibit a cubic-ice structure, with the (110) plane aligned with the (110) K-feldspar surface. Our results shed new light on the atomistic mechanism of this process and provide an explanation for the high ice nucleation efficiency of K-feldspar.
@@ Conclusions
Our MD simulations show that the (110) surface of K-feldspar is responsible for its high ice nucleation efficiency. A comprehensive analysis of alternative feldspar surfaces shows no evidence of enhanced ice nucleation. Previous work had suggested that the high-energy feldspar (100) surface, exposed at surface defects, provides preferential sites for ice nucleation, yet our findings are inconsistent with that hypothesis. Our simulations also provide several key microscopic insights into this phenomenon. First, we find that the high ice nucleation efficiency of the (110) feldspar surface is explained by its unique ability to structure liquid water into an arrangement closely resembling that on the (110) surface of cubic ice, thereby providing an optimal template for ice nucleation and growth. Furthermore, the ice clusters formed over the (110) feldspar surface are approximately hemispherical, with minor faceting, and mostly composed of cubic ice, a metastable polymorph with respect to hexagonal ice. The picture emerging from our study is that atmospheric ice nucleation in mixed-phase clouds occurs at feldspar defects where patches of the (110) surface are exposed. The initial nuclei adopt a cubic ice structure but subsequently transform into hexagonal ice as they grow to macroscopic dimensions. Our work provides a clear atomic-level explanation for the exceptionally high ice nucleation efficiency of K-feldspar, which has been shown to be the dominant ice nucleating particle in the atmosphere.

### v2

# Molecular mechanism of heterogeneous ice nucleation in the atmosphere

@@ Abstract
Mineral dust aerosols strongly influence Earth’s climate by acting as ice-nucleating particles (INPs). Feldspar minerals, particularly K-feldspar, are recognized as dominant INPs, and a previous study attributed this behavior to (100) surfaces exposed at defects. Using machine-learning molecular dynamics simulations, we systematically investigate ice nucleation on all K-feldspar surfaces. We identify the (110) surface, exposed at defects such as steps, as the most active plane for ice formation. This surface uniquely structures interfacial water into an arrangement resembling that on the (110) surface of cubic ice, providing an optimal template for nucleation. Using advanced sampling, we directly observe the formation of clusters with cubic-ice structure and their orientation agrees with experiment. These results provide a molecular-level explanation of how ice forms in our planet’s atmosphere.
@@ Introduction
Figure 1: K-feldspar serves as an effective ice-nucleating particle in the atmosphere owing to structural defects that promote ice nucleation. (A) Schematic illustration of the heterogeneous ice nucleation at K-feldspar defects with exposed non-perfect cleavage planes. The pink and green dashed lines highlight the first and second prismatic surfaces of ice, which are approximately parallel to the (100) and (110) planes of feldspar, respectively. (B) The triclinic crystal structure of microcline K-feldspar. (C) Crystallographic planes of K-feldspar: (001), (010), (100), (110), ($\bar{1}$10), and ($\bar{2}$01). (D) Configuration for molecular dynamics simulations with one termination of the K-feldspar (110) surface in contact with water above.
In this work, we use MD simulations to study a broad range of K-feldspar surfaces exposed to water, including the (001), (010), (100), (110), ($\bar{1}$10), and ($\bar{2}$01) surfaces. The interatomic forces driving these simulations are derived from a machine-learning interatomic potential (MLP) trained to achieve a very high accuracy using quantum-mechanical electronic-structure theory calculations. Our simulations reveal that the K-feldspar (110) surface structures interfacial liquid water into an arrangement closely resembling that on the (110) surface of cubic ice (equivalent to the secondary prismatic surface of hexagonal ice). We also carried out enhanced sampling molecular simulations in order to bring ice nucleation, which is a rare event in the timescale of standard MD simulations, within reach. The ice clusters that nucleate at the (110) surface during these simulations exhibit a cubic-ice structure, with the (110) plane aligned with the (110) K-feldspar surface. In contrast, none of the other feldspar surfaces show signatures of high ice nucleation potency. Our results shed new light on the molecular mechanism for heterogeneous ice nucleation on K-feldspar, which is the main route for ice formation in the atmosphere.
@@ Conclusions
Our MD simulations show that the (110) surface of K-feldspar is responsible for its high ice nucleation efficiency. A comprehensive analysis of alternative feldspar surfaces shows no evidence of enhanced ice nucleation. Our results also provide two key molecular-level insights into this phenomenon. First, we find that the high ice nucleation efficiency of the (110) feldspar surface is explained by its unique ability to structure liquid water into an arrangement closely resembling that on the (110) surface of cubic ice, thereby providing an optimal template for ice nucleation and growth. Second, the ice clusters formed over the (110) feldspar surface are approximately hemispherical, with minor faceting, and mostly composed of cubic ice, a metastable polymorph with respect to hexagonal ice. These findings stand in contrast to previous work, which had suggested that the high-energy feldspar (100) surface provides preferential sites for ice nucleation. Our results contradict this hypothesis and instead support a new interpretation in which atmospheric heterogeneous ice nucleation occurs at feldspar defects exposing patches of the (110) surface. In this new picture, the initial nuclei adopt a cubic ice structure but subsequently transform into hexagonal ice as they grow to macroscopic dimensions. Our work provides a clear molecular-level explanation for the exceptionally high ice nucleation efficiency of K-feldspar, which has been shown to be the dominant ice nucleating particle in the atmosphere.